OpenStudy (anonymous):

evaluate 6(cos80deg + i sin 80deg)/3(cos35deg + i sin35deg)

OpenStudy (anonymous):

Do u kne Eucler Formula....Use it....

OpenStudy (anonymous):

we didn't takle that in the school :(

OpenStudy (anonymous):

cos80deg + i sin 80deg = e^ (i 80) cos35deg + i sin35deg = e^ (i 35)

OpenStudy (anonymous):

6e^ (i 80)/3 e^ (i 35)

OpenStudy (anonymous):

2 e^ (i 45) = 2 ( cos 45 + i sin45)........nw just plug value for cos and sin.... hope this helps

OpenStudy (anonymous):

it helps a lot my friend :))

OpenStudy (anonymous):

the choices are a)sqroot of 2 (1 + i) b) sqroot of 2 (1-i) c) 2(1+i) d) 2(1-i)